If 3x + (y-2)i = (5-2x) + (3y-8)i, then (x, y) is?
Another one: If (x + iy)2 = 3+4i, then (x, y) is?
How am I supposed to proceed?

Answer by blueplanet(7)

Both equations can be solved by equating their real parts and by equating their imaginary parts.
To solve 3x + (y-2)i = (5-2x) + (3y-8)i, we can approach the problem by equating the real parts first:
3x = 5-2x
=> x=1
Then, equating the imaginary parts, we get:
y-2 = 3y-8
y = 3
So, (x,y) = (1,3).
You can use a similar approach to resolve (x + iy)2 = 3+4i. If my calculation is correct, the answer should be (x,y) = (3/2, 2).